Solve. –9(x – 3)^2 = –7
Another difficult problem that I need help with, please.

Answer by checkley75(3666)

-9(X-3)^2=-7 POWERS FIRST
-9(X^2-6X+9)=-7 NEXT MULTIPLICATION
-9X^2+54X-81=-7 NOW COMBINE LIKE TERMS
-9X^2+54X-81+7=0 NOW MULTIPLY BY -1 WE GET
9X^2-54X+74=0 NOW FACTOR USING THE QUADRATIC EQUATION
X=(-B+-SQRT[B^2-4AC])/2A
X=(54+-SQRT[2916-4*9*74])/2*9
X=(54+-SQRT[2916-2664])/18
X=(54+-SQRT[252])/18
X=(54+-15.87)/18
X=(54+15.87)/18
X=69.87/18
X=3.88 ANSWER
X=(54-15.87)/18
X=38.13/18
X=2.12 ANSWER